find the value of x;
 
2x-1=5
 
plixx find it fast  and do it now

Grade:12th pass

1 Answers

Harshit Singh
3 years ago
Welcome to AskIITians

2x-1=5

2x= 5+1
2x=6
x=6/2
x=3

Thanks